40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4670     Entry time: Mon May 9 17:23:25 2011
Author: Suresh 
Type: Update 
Category: RF System 
Subject: RF Cables near LSC Rack 

[Steve, Suresh]

We started to clean up the RF cables (heliax and PD interface cables)  at the LSC rack.

We have pulled out all the RF cables from the small hole on the side-board close to floor.  Passing the cables through this hole makes some of the cables much too short for good strain relief.  So we removed the side panel on the vacuum tube side and are going to pass the cables into the rack from there at about waist height.  We now have plenty of cable lengths to tie them off to the rack at several points.

We have traced all the available Heliax cables and have attached blank tags to them.  We have allocated some cables to REFL11, REFL55 and AS55.  These are therefore back in working order.  We have also taken stock of the available PD interface cables.  They do not have consistent names on both ends of the cable and we will identify and label the ends tomorrow.

MC is locked.  The auto-locker works fine.

Handing over the system for night time interferometer work now.  Will continue with the cabling tomorrow.

 

ELOG V3.1.3-